OpenStudy (anonymous):

r1=1,4,7 r2=2,5,8 r3=3,6,9 (vectors) Linear algebra says that these vectors must also lie in a plane. there must be many combinations with y1r1+y2r2+y3r3=0 find two sets of y's please help!!

OpenStudy (anonymous):

For linear dependence \[a(1,4,7)+b(2,5,8)+c(3,6,9)=(0,0,0)\] For any non zero linear combination of a,b,c. Solving we get, \[a+2b+3c=0\] \[4a+5b+6c=0\] \[7a+8b+9c=0\] Subtracting 3rd from 2nd \[3(a+b+c)=0\] \[b+c=-a\] from 1st \[a=-2b-3c\] \[2b+3c=b+c\] \[b=-2c\] \[a=-b-c=-b+2c=c\] Thus linear combination \[c(1,-2,1)\] for all c in real numbers

OpenStudy (anonymous):

thank you a lot!!!